# AutoRAN: Automated and Zero-Touch Open RAN Systems

### v2

@@ III-E High-Level Intents to Represent Network Status
Therefore, we design AutoRAN with a mechanism to provide high-level input for the system configuration, which is then translated into actionable templates, pipeline triggers, and inputs that are automatically applied to the system. As shown in the top left part of Figure 2, users can express requests to instantiate services or tests at a high level (e.g., “deploy a 5G base station”, which are parsed through an LLM that matches the intent with specific templates to be applied to the system. We provide details on the implementation of this mechanism in Section V.
@@ III-F Observability
Finally, once the system is configured, deployed, and operational, the AutoRAN design also embeds frameworks and procedures to automatically monitor the system status and coordinate with the orchestrator to maintain the desired system state. AutoRAN leverages Prometheus, OpenShift APIs, and Observium [27] to track application performance, resource consumption, and network status in real time. These tools work together to capture and visualize key metrics—such as CPU load, memory usage, and network traffic, enabling proactive identification of anomalies, informed resource management, and continuous improvements to overall system reliability and efficiency.
@@ IV AutoRAN System Components
In this section, we will describe the core components that enable the automation performed by AutoRAN. Section IV-A describes the OpenShift cluster leveraged by AutoRAN, with operators and configuration profiles detailed in Sections IV-B and IV-C, respectively. Section IV-D describes how to integrate and deploy RAN workloads using the O-RAN 7.2 split, with and without a layer-1 accelerator, and with USRP radios based on a 8.1 split. Finally, Section IV-E details how to manage and build images for heterogeneous architectures, while Section IV-F provides insights on additional software components of the cluster.
@@ IV-A The Automation Cluster
Compute nodes are differentiated into control-plane and worker nodes. The former are used to host automation, control and monitoring functionalities of the cluster, and core/RIC/AI workloads. The latter nodes are optimized to host RAN workloads and services via low-latency configurations and GPU acceleration capabilities.
Servers connect to the cluster via $10$ Gbps connections to the backhaul switch, and to the fronthaul interface and RUs via dedicated $100$ Gbps connections. AutoRAN includes both nodes with x86 and ARM architectures. The ARM nodes, i.e., Grace Hopper MGX GH200, are used to host the latest release of the high-performance NVIDIA ARC solution.
Finally, AutoRAN also includes a persistent storage via a dedicated NAS. We separate computing from storage resources for resilience, thus obtaining a stateless cluster. Specifically, the NAS used by AutoRAN includes $12$ TB of storage in RAID 3 using the open-source TrueNAS solution [28] to share volumes over the NFS protocol. The RAID 3 redundancy level implements a replica of the data over three separate disks, where two are used for data, while the third one is used for parity, which makes the usable disk space $8$ TB.
@@ IV-B Operators for Automatic Configuration
PTP Operator. RAN systems require precise timing synchronization across their components. In general, clock synchronization is achieved through the Network Time Protocol (NTP), which is able to provide ms accuracy over Local Area Network (LAN) setups. However, the level of accuracy delivered by NTP is not sufficient to support RAN applications, where the Open Fronthaul interface requires nanosecond accuracy. We therefore use PTP, a protocol that shares a synchronization signal over Ethernet, to sync all nodes and radios to a common clock source—a Qulsar clock with a GPS input in our case. The Linux implementation of the protocol uses two pieces of software, namely ptp4l and phc2sys, which are installed via the OpenShift PTP operator. The operator also gives the possibility of increasing the priority of these processes, allowing one to use them on shared CPUs rather than on a dedicated CPU per process, saving one extra core for other workloads. Listing 1 shows the PTP configuration adapted for the Grace Hopper server.
@@
SR-IOV Operator. This operator is in charge of creating virtual slices—or virtual functions—of the physical NIC interfaces and to expose them as objects inside OpenShift. In OpenShift and Kubernetes deployments, these are created using configuration files called SR-IOV Networks Node Attachment. After splitting the interfaces, OpenShift implements Network attachments to specify the IP addressing of the network and the VLAN tag. As it will be discussed in Section IV-D, NVIDIA ARC deployments require interfaces to be passed as untagged whereas OAI requires the tag to be set on the fronthaul port. This becomes key as different deployments need different settings on the physical port, and it enables parallelization of multiple workloads, and thus concurrent accelerated workloads on the same compute nodes. Common NVIDIA ARC deployments are designed to directly use physical network interfaces, whereas our SR-IOV approach, which enables running the L1 over a virtualized interface, makes it possible to parallelize multiple throughput-demanding applications. To the best of our knowledge, we are the first in following this approach with NVIDIA ARC.
@@ IV-C Configuration Profiles
Due to the stringent latency and processing requirements of RAN workloads, it is important to make sure that host machines use an appropriate kernel version and are configured to meet such requirements. This includes setting isolated CPUs, HugePages and disabling energy saving functionalities that might put the processor in an idle state and decrease performance. As mentioned in Section IV-A, we use labels to automatically apply configurations to nodes. Nodes with the same physical hardware and configurations are grouped together in an Machine Configuration Pool (MCP). Then, for each MCP, we apply the same set of configurations.
@@
Profiles linked to a label are immediately applied to each new labeled node. Nodes update in a rolling strategy and, at the end of the update, their configurations are aligned, which eliminates the risk of stale drivers. This flexibility in node labeling and zero-touch configuration is even more relevant for extendability and scalability of the framework. Finally, it is worth mentioning that differently from bare-metal or Kubernetes-based deployments, OpenShift only supports generic or real-time kernels (from no other sources than Red Hat itself), and does not support the low-latency ones generally recommended for gNB protocol stacks. Since at this time the NVIDIA L1 deployment requires drivers that are not available for Red Hat’s real-time kernel, we adopt the generic Linux kernel on all the cluster nodes. Experimental evidence in [13] also shows the instabilities of real-time kernels for RAN deployments, which results in poorer performance when compared to the generic one.
@@ IV-D Integrating RAN Deployments
Figure 4: ARC deployment in AutoRAN.
@@
Figure 6: End-to-end automated testing diagram.
@@ IV-E Container Image Building
To maximize image reuse and reduce build times, images are built in a chained manner, as shown in Figure 5. Specifically, different gNB images (e.g., OAI, srsRAN) and their subsequent versions (e.g., among different weekly tags in the case of OAI) might share common dependencies, e.g., in terms of NIC drivers, DPDK, etc. Instead of rebuilding all the dependencies for each new image, we maintain dedicated images with the required dependencies installed but no gNB software. We, then, use these images as a base to build the gNB images, thus speeding up the build process at the cost of additional images (i.e., the dependencies images) stored on the registry.
@@ IV-F Additional Software Components
In addition to the RAN workloads, a fully operational Open RAN requires several other elements. The core network is in charge of operations such as user authentication and the creation of Packet Data Unit (PDU) sessions. Among the various core networks available in the literature, we deployed Open5GS.
Another key element of an O-RAN deployment is the RIC, a component that enables observability and control of the network [30]. Specifically, Near-real-time (RT) control is achieved via applications called xApps deployed on the Near-RT RIC, while Non-RT control via rApps deployed on the Non-RT RIC We deploy an OSC Near-RT RIC (“E” release) and connect it to OAI via the E2 agent provided in [8, 31] and publicly available as part of the OpenRAN Gym framework [32]. Finally, the cluster also hosts a non-RT RIC to enable observability at larger time scales.
@@ V-A Intent-based Instantiation and Testing
In the following paragraphs, we describe how we leverage AutoRAN to simplify and streamline network instantiation procedures and testing via intent-based deployment and testing, leveraging natural language processing. The goal of this feature is to eliminate complexity related to network configuration and allow even non-expert users to deploy and test a full-fledged network using only Natural Language (NL). Examples of possible intents include queries such as “deploy a 5G gNB with OAI and NVIDIA Aerial Research Cloud (ARC) or “perform a 15 Mbps iPerf test with 3 UEs. In order to process NL, we opt for using LLMs as they provide capabilities to parse text, understand context and need, and convert them into structured and machine-understandable directives. Given the large availability of pre-trained models with excellent performance in accomplishing a variety of tasks under different benchmarks, we opt for using a pre-trained LLM (Qwen 2.5) in an agentic setup, i.e., where the LLM makes calls to a set of tools (defined thereafter) in iterative loops.
@@
  • First, it should be noted that not all the RAN components available in AutoRAN are compatible with each other. For example, OAI supports the ARC L1, while srsRAN does not, meaning we need to prevent the LLM from trying to configure a srsRAN base station with ARC, which would result in a non-working deployment. It is therefore fundamental to ensure that the LLM only selects compatible components and configurations (e.g., OAI with ARC, plain OAI/srsRAN, as shown in Table III). For this reason, we store a compatibility graph, indicating which network components (DU, CU, RU, Core Network) are compatible with one another. Similarly, for each test type, a graph of the list of compatible parameters is fed to the LLM (for example, tests using different traffic generators, such as Multi-Generator (MGEN) can include parameters such as the distribution of traffic, while iPerf tests do not). In the workflow of Figure 7, this graph is consumed when validating values.
@@ V-B Deployment Workflow
Each pipeline consists of multiple tasks that need to be executed to successfully run the pipeline. These tasks are associated to operations that span from sending the intent of the user to the LLM, to converting the output of the latter into a deployment object specialized to the physical infrastructure, to configuring and instantiating the gNB components (e.g., CUs and DUs) and connect them to the radio devices used for the test (e.g., either USRP SDRs or commercial RUs) and to the core network. After instantiating the specified components in the form of virtualized containers, the gNB software is started via a call to Flask APIs—or multiple calls in the case of gNB accelerated by the NVIDIA ARC solution, i.e., one for the CU/DU-high container and one for the DU-low container with CUDA Baseband (cuBB)—that they expose.
@@ V-C Testing Workflow
For each UE, this database—automatically updated by the UEs via cron jobs—records information such as UE hostname, International Mobile Subscriber Identity (IMSI), IP address, location, and model and serial number of the mini-PC controlling the Sierra Wireless 5G modem. Finally, once UEs are connected to the gNB, the specified traffic will start (for instance using tools such as iPerf), and results of the test will be collected and analyzed. In our tests, we deploy multiple iPerf servers to generate the traffic that the UEs request during the test, and each instance is deployed as a pod. Specific parameters of the test, such as server hostname and port, can be specified in the test files.
After the specified tests have been executed, performance metrics are extracted and stored in an InfluxDB database—together with test and protocol stack logs—and visualized via a Grafana dashboard, which also lets users compare the latest test results to historical data.
This, together with the automated update of the gNB protocol stack is key for observing how new releases of the protocol stack software affect the network performance, enabling users to quickly spot degradations with respect to previous tests.
@@ VI Experimental Evaluation
The section discusses several results that profile the performance and effectiveness of the proposed AutoRAN approach. Specifically, in Section VI-A we review the time needed to onboard a new node in the cluster, comparing automation and manual solutions. In Section VI-B we describe the LLM deployment and report related metrics. In Section VI-C, we analyze the time needed to execute deployment and testing workflows using AutoRAN, showing how a network based on microservices can be instantiated in a matter of seconds. In Section VI-D, we provide performance metrics obtained from the deployed gNBs. Finally, in Section VI-E we analyze coexistence between RAN and other generic workloads.
@@ VI-A Onboarding Nodes on the AutoRAN Cluster
AutoRAN is designed to ensure scalability. Even the extension of the cluster follows the same principle. To showcase the benefits of automated node creation and configuration, in this section we provide a comparison between adding a new RAN worker node manually and using AutoRAN. Manually deploying a NVIDIA ARC node requires performing several error-prone steps that need to be repeated on each compute node of the cluster just to set up the basic infrastructure. This usually requires additional effort to maintain the node with up-to-date drivers and software.
Figure 8: Steps required to provision a new node.
Using the profiles and configurations offered by AutoRAN, extending the cluster with an already configured profile requires minimal human intervention, as shown in Figure 8 (i.e., it is only required to add the node to the cluster—using coreos-installer—and assign a label to it—using oc label), and guarantees a completely configured node ready to accept any targeted workload in approximately 40 minutes—without the same level of effort of manually configuring a node. After their initial provisioning, nodes can be repurposed for different roles (e.g., edge vs RAN node) by modifying their labels. After this operation, nodes will reboot and auto-configure themselves as required.

### v3

@@ III-E High-Level Intents to Represent Network Status
Therefore, we design AutoRAN with a mechanism to provide high-level input for the system configuration, which is then translated into actionable templates and pipeline triggers that are automatically applied to the system. As shown in the top left part of Figure 2, users can express high-level requests to instantiate services or tests (e.g., “deploy a 5G base station”, which are parsed through an LLM matching the intent with templates to be applied to the system. We provide details on the implementation of this mechanism in Section V.
@@ III-F Observability
Finally, once the system is configured, deployed, and operational, the AutoRAN design also embeds procedures to automatically monitor the system status and coordinate with the orchestrator to maintain the desired system state. AutoRAN leverages Prometheus [28], OpenShift APIs, and Observium [29] to track application performance, resource consumption, and network status in real time. These tools work together to capture and visualize key metrics, e.g., CPU load, memory usage, and network traffic, and enable identification of anomalies, resource management, and continuous improvements to the overall system reliability and efficiency.
@@ IV AutoRAN System Components
In this section, we will describe the core components that enable the automation performed by AutoRAN. Section IV-A describes the OpenShift cluster leveraged by AutoRAN, with operators and configuration profiles detailed in Sections IV-B and IV-C, respectively. Section IV-D describes how to integrate and deploy RAN workloads using the O-RAN 7.2 split, with and without a layer-1 accelerator, and with SDRs based on a 8.1 split. Finally, Section IV-E details how to manage and build images for heterogeneous architectures, while Section IV-F reviews additional software components of the cluster.
@@ IV-A The Automation Cluster
Compute nodes are differentiated into control-plane and worker nodes. The former are used to host automation, control and monitoring functionalities of the cluster, as well as generic core network, RIC, and AI workloads. The latter are optimized to host RAN workloads and services via low-latency configurations and GPU acceleration capabilities. However, concurrently supporting RAN and generic cloud computing workloads comes with specific requirements on the underlying infrastructure. As detailed in Table I, we select high-performance servers with different architectures (x86 and ARM) and acceleration capabilities (e.g., the Grace Hopper MGX GH200 ARM nodes used to run the GPU-accelerated NVIDIA ARC solution) to meet the stringent RAN requirements. Key considerations also include maximizing physical CPU cores, as RAN workloads benefit from disabled hyperthreading functionalities [30], and require isolated cores for optimal performance.
Servers connect to the cluster via $10$ Gbps connections to the backhaul switch, and to the fronthaul interface and RUs via dedicated $100$ Gbps connections to satisfy fronthaul traffic requirements (e.g., each $4\!\times\!4$ MIMO RU with standard 9-bit compression requires a $10$ Gbps connection). AutoRAN also consolidates multiple DUs on each node, thus calling for a $100$ Gbps NIC. A PTP-compatible NIC is also needed for carrying fronthaul traffic while maintaining precise synchronization through PTP timestamping. Finally, AutoRAN also includes a persistent storage via a dedicated NAS. We separate computing from storage resources for resilience, thus obtaining a stateless cluster. Specifically, the NAS used by AutoRAN includes $12$ TB of storage in RAID 3 using the open-source TrueNAS solution [31] to share volumes over the NFS protocol. The RAID 3 redundancy level implements a replica of the data over three separate disks, where two are used for data, while the third one is used for parity, which makes the usable disk space $8$ TB.
@@ IV-B Operators for Automatic Configuration
PTP Operator. RAN systems require precise timing synchronization across their components. In general, clock synchronization is achieved through the Network Time Protocol (NTP), which is able to provide millisecond accuracy over Local Area Network (LAN) setups. However, the level of accuracy delivered by NTP is not sufficient to support RAN applications, where the Open Fronthaul interface requires nanosecond accuracy. We therefore use PTP, a protocol that shares a synchronization signal over Ethernet, to sync all nodes and radios to a common clock source—a Qulsar clock with a GPS input in our case. The Linux implementation of the protocol uses two pieces of software, namely ptp4l and phc2sys, which are installed via the OpenShift PTP operator. The operator also gives the possibility of increasing the priority of these processes, allowing one to use them on shared CPUs rather than on a dedicated CPU per process, saving one extra core for other workloads. Listing 1 shows the PTP configuration adapted for the Grace Hopper server.
@@
SR-IOV Operator. This operator is in charge of creating virtual slices—or virtual functions—of the physical NIC interfaces and to expose them as objects inside OpenShift. In OpenShift and Kubernetes deployments, these are created using configuration files called SR-IOV Networks Node Attachment. After splitting the interfaces, OpenShift implements Network attachments to specify the IP addressing of the network and the VLAN tag. As it will be discussed in Section IV-D, NVIDIA ARC deployments require interfaces to be passed as untagged, whereas OAI requires the tag to be set on the fronthaul port. This becomes key as different deployments need different settings on the physical port, and it enables parallelization of multiple workloads on the same compute nodes. Additionally, without using SR-IOV, a NIC can only be accessed by a single container (e.g., a DU) at a time, whereas, with SR-IOV, it is possible to run multiple applications requiring access to same physical interface by connecting each of them to the virtual interfaces created by SR-IOV. This is used, for instance to parallelize multiple NVIDIA ARC deployments by running the L1 over the virtualized SR-IOV interface, where we partition each NIC into 8 virtual interfaces. For this number of partitions, we do not observe a significant increase in latency compared to bare metal or Docker-based setups. To the best of our knowledge, we are the first in following this approach with NVIDIA ARC.
@@ IV-C Configuration Profiles
Due to the stringent latency and processing requirements of RAN workloads, it is important to make sure that host machines use an appropriate kernel version and are configured to meet such requirements. This includes setting isolated CPUs, HugePages and disabling energy saving functionalities that might put the processor in an idle state and decrease performance. We use labels to automatically apply configurations to nodes upon their addition to the cluster. Nodes with the same physical hardware and configurations are grouped together in an Machine Configuration Pool (MCP). Then, for each MCP, we apply the same set of configurations.
@@
Profiles linked to a label are immediately applied to each new labeled node. Nodes update in a rolling strategy and, at the end of the update, their configurations are aligned, which eliminates the risk of stale drivers. This flexibility in node labeling and zero-touch configuration is even more relevant for extendability and scalability of the framework. Indeed, multiple profiles can be prepared in advance and applied to MCPs. If a node—or a set of nodes—needs to be adapted for different workloads (from RAN to generic cloud computing or high-performance computing), it is only necessary to relabel the node so as to make it part of a different MCP with different configurations, and the node gets automatically provisioned.
Finally, it is worth mentioning that differently from bare-metal or Kubernetes-based deployments, OpenShift only supports generic or real-time kernels (from no other sources than Red Hat itself), and does not support the low-latency ones generally recommended for gNB protocol stacks. Since at this time the NVIDIA L1 deployment requires drivers that are not available for Red Hat’s real-time kernel, we adopt the generic Linux kernel on all the cluster nodes. Experimental evidence in [14] also shows the instabilities of real-time kernels for RAN deployments, which results in poorer performance when compared to the generic one.
@@ IV-D Integrating RAN Deployments
All gNBs based on the 7.2 split and without GPU accelerator share the same deployment structure. Therefore, when a new radio—real or emulated as RuSIM—gets integrated into AutoRAN, the only modification required is in the template of the configuration file. The same applies, for both L1 and L2 configuration files, to the NVIDIA ARC deployment.
@@
Figure 6: E2E automated testing diagram.
@@ IV-E Container Image Building
To maximize image reuse and reduce build times, images are built in a chained manner, as shown in Figure 5. Specifically, different gNB images (e.g., OAI, srsRAN) and their subsequent versions (e.g., among different weekly tags in the case of OAI) might share common dependencies, e.g., in terms of NIC drivers, DPDK, etc. Instead of rebuilding all the dependencies for each new image, we maintain dedicated images with the required dependencies installed but no gNB software. We, then, use these images as a base to build the gNB images, thus speeding up the build process at the cost of additional images (i.e., the dependencies images) stored on the registry. Since AutoRAN is deployed across heterogeneous server architectures (ARM and x86) and CPU models (e.g., Cascade Lake, Sapphire Rapids), we adopt an MCP-specific build approach. To ensure that each RAN image fully leverages the capabilities of the hosting node, images are built directly on the nodes where the corresponding workloads will be executed.
@@ IV-F Additional Software Components
In addition to the RAN workloads, a fully operational Open RAN requires several other elements. The core network is in charge of operations such as user authentication and the creation of Packet Data Unit (PDU) sessions. Among the various core networks available in the literature, we deployed Open5GS. However, AutoRAN is not bound to a specific core network, and we decided to focus on Open5GS due to of its high performance and ease of installations in Kubernetes-based deployments, where we deploy multiple independent replicas with different PLMNs for different tenants.
Another key element of an O-RAN deployment is the RIC, a component that enables observability and control of the network [33]. Specifically, Near-Real-Time (Near-RT) control is achieved via applications called xApps deployed on the Near-RT RIC, while Non-RT control via rApps deployed on the Non-RT RIC We deploy an OSC Near-RT RIC (“E” release) and connect it to OAI via the E2 agent provided in [9, 34] and publicly available as part of the OpenRAN Gym framework [35]. Finally, the cluster also hosts a non-RT RIC to enable observability at larger time scales.
@@ V-A Intent-based Instantiation and Testing
In the following paragraphs, we describe how we leverage AutoRAN to simplify and streamline network instantiation procedures and testing via intent-based deployment and testing, leveraging natural language processing. The goal of this feature is to eliminate complexity related to network configuration and allow even non-expert users to deploy and test a full-fledged network using only Natural Language (NL). Examples of possible intents include queries such as “deploy a 5G gNB with OAI and NVIDIA Aerial Research Cloud (ARC) or “perform a 15 Mbps iPerf test with 3 UEs. In order to process NL, we opt for using LLMs as they provide capabilities to parse text, understand context and need, and convert them into structured and machine-understandable directives. Given the large availability of pre-trained models with excellent performance for a variety of tasks under different benchmarks, we opt for using a pre-trained LLM (Qwen 2.5) in an agentic setup, i.e., where the LLM makes calls to a set of tools (defined thereafter) in iterative loops.
@@
  • First, it should be noted that not all the RAN components available in AutoRAN are compatible with each other. For example, OAI supports the ARC L1, while srsRAN does not, meaning we need to prevent the LLM from trying to configure a srsRAN base station with ARC, which would result in a non-working deployment. It is therefore fundamental to ensure that the LLM only selects compatible components and configurations (e.g., OAI with ARC, plain OAI/srsRAN, as shown in Table III). For this reason, we store a compatibility graph, indicating which network components (DU, CU, RU, Core Network) are compatible with one another. Similarly, for each test type, a graph of the list of compatible parameters is fed to the LLM (for example, tests using different traffic generators, such as Multi-Generator (MGEN) can include parameters such as the distribution of traffic, while iPerf tests do not). In the workflow of Figure 7, this graph is consumed when validating values. When a new technology (CU, DU, RU) becomes available and integrated, the dependability graph is updated to reflect the compatibility between that specific element and the others available in AutoRAN.
@@ V-B Deployment Workflow
Each pipeline consists of multiple tasks that need to be executed to successfully run the pipeline. These tasks are associated to operations that span from sending the intent of the user to the LLM, to converting the output of the latter into a deployment object specialized to the physical infrastructure, to configuring and instantiating the gNB components (e.g., CUs and DUs) and connect them to the radio devices used for the test (e.g., either USRP SDRs or commercial RUs) and to the core network. After instantiating the specified components in the form of virtualized containers, the gNB software is started via a call to Flask APIs—or multiple calls in the case of gNB accelerated by the NVIDIA ARC solution, i.e., one for the CU/DU-high container and one for the DU-low container with CUDA Baseband (cuBB)—that they expose. Thus, by leveraging OpenShift’s built-in virtualization and parallelization capabilities, AutoRAN can execute multiple experiments concurrently, provided sufficient hardware resources are available and there is no interference between the tested gNBs.
@@ V-C Testing Workflow
For each UE, this database—automatically updated by the UEs via cron jobs—records information such as UE hostname, International Mobile Subscriber Identity (IMSI), IP address, location, and model and serial number of the mini-PC controlling the Sierra Wireless 5G modem. This automation enables continuous scalability of the system without requiring manual updates to the testing framework or UE database. Finally, once UEs are connected to the gNB, the specified traffic will start (for instance using tools such as iPerf), and results of the test will be collected and analyzed. In our tests, we deploy multiple iPerf servers to generate the traffic that the UEs request during the test, and each instance is deployed as a pod. Specific parameters of the test, such as server hostname and port, can be specified in the test files. After the specified tests have been executed, performance metrics are extracted and stored in an InfluxDB database—together with test and protocol stack logs—and visualized via a Grafana dashboard, which also lets users compare the latest test results to historical data. This, together with the automated update of the gNB protocol stack is key for observing how new releases of the protocol stack software affect the network performance, enabling users to quickly spot degradations with respect to previous tests.
Figure 8: Steps required to provision a new node.
@@ VI Experimental Evaluation
The section discusses several results that profile the performance and effectiveness of the proposed AutoRAN approach. Specifically, in Section VI-A we review the time needed to onboard a new node in the cluster, comparing automation and manual solutions. In Section VI-B we describe the LLM deployment and report related metrics. In Section VI-C, we analyze the time needed to execute deployment and testing workflows using AutoRAN, showing how a network based on microservices can be instantiated in a matter of seconds. In Section VI-D, we provide performance metrics obtained from the deployed gNBs. In Section VI-E we analyze coexistence between RAN and other generic workloads. Finally, in Section VI-F, we provide insights on AutoRAN’s resiliency.
@@ VI-A Onboarding Nodes on the AutoRAN Cluster
AutoRAN is designed to ensure scalability. Even the extension of the cluster follows the same principle. To showcase the benefits of automated node creation and configuration, in this section we provide a comparison between adding a new RAN worker node manually and using AutoRAN. Manually deploying a NVIDIA ARC node requires performing several error-prone steps that need to be repeated on each compute node of the cluster just to set up the basic infrastructure. This usually requires additional effort to maintain the node with up-to-date drivers and software. Using the profiles and configurations offered by AutoRAN, extending the cluster with an already configured profile requires minimal human intervention, as shown in Figure 8 (i.e., it is only required to add the node to the cluster—using coreos-installer—and assign a label to it—using oc label), and guarantees a completely configured node ready to accept any targeted workload in approximately 40 minutes—without the same level of effort of manually configuring a node. After their initial provisioning, nodes can be repurposed for different roles (e.g., edge vs RAN node) by modifying their labels. After this operation, nodes will reboot and auto-configure themselves as required.